OpenStudy (alexh107):

Trig equations help: Find all solutions in the interval [0, 2π). cos x = sin x Really in need of help with this concept so if anyone can explain it to me it would be very appreciated.

OpenStudy (anonymous):

look on the unit circle to see where the first and second coordinates are the same you got a unit circle cheat sheet?

OpenStudy (anonymous):

cosine is the first coordinate, sine is the second there are two points on the unit circle where the first and second coordinates are the same it is on the last page1

OpenStudy (alexh107):

Okay one second I'm looking at it now.

OpenStudy (alexh107):

I'm not really sure which two points on the circle have the same first and second coordinates. Each quadrant make a different coordinate negative/positive so isn't that impossible?

jimthompson5910 (jim_thompson5910):

hint: look in quadrant 1 and quadrant 3

OpenStudy (alexh107):

Is it pi/4 and 5pi/4? I feel like it's really obvious but I'm just not seeing it.

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (alexh107):

Would that be the final answer then?

jimthompson5910 (jim_thompson5910):

you just said them both: pi/4 and 5pi/4

jimthompson5910 (jim_thompson5910):

cos(pi/4) = sqrt(2)/2 sin(pi/4) = sqrt(2)/2 this is shown at the point in Q1 with both coordinates equal to sqrt(2)/2

jimthompson5910 (jim_thompson5910):

cos(5pi/4) = -sqrt(2)/2 sin(5pi/4) = -sqrt(2)/2 this is shown at the point in Q3 with both coordinates equal to -sqrt(2)/2

OpenStudy (alexh107):

Oh okay, thank you so much!

jimthompson5910 (jim_thompson5910):

I guess another way to do it is to square both sides and use an identity sin(x) = cos(x) sin^2(x) = cos^2(x) sin^2(x) = 1-sin^2(x) sin^2(x)+sin^2(x) = 1 2*sin^2(x) = 1 keep going to solve for x You'll have to check the possible solutions (some possible solutions will be extraneous)

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (alexh107):

That makes a lot of sense too, thank you!